# Dynamic Conditional SKEPTIC

### v1

@@ 1 Introduction
Multivariate GARCH models analyze the dependence structure of a vector of asset returns through the dynamic estimation of its covariance matrix. When the number of assets is large, these models face computational issues in the estimation procedure that returns poor performance due to the curse of dimensionality. To this end, the recent works of engle2019large and de2022large robustify the Dynamic Conditional Correlation (DCC) (engle2002dynamic), one of the most commonly used in the class of multivariate GARCH models, for estimating time-varying correlations in financial assets, including a nonlinear shrinkage (ledoit2015spectrum) that improves the estimation of the unconditional correlation matrix and relying on the composite likelihood method (pakel2021fitting) that provides computationally feasible estimation of the DCC model in large-dimensional settings.
However, a limitation of the DCC model remains the assumption that asset returns follow a normal distribution. This assumption does not adequately capture the empirical stylized facts (cont2001empirical) which are commonly observed in asset return distributions (bai2003kurtosis), particularly the skewness and leptokurtosis. To formalize the concept, let $\boldsymbol{X}_{t}=(X_{1,t},\ldots,X_{p,t})^{\prime}$ represent the $p$-dimensional random vector of asset returns, the main assumption is that $\boldsymbol{X}_{t}\sim\operatorname{N}(0,H_{t})$, where $H_{t}$ is the covariance matrix. The difference between the normal distribution and the empirical distribution of $\boldsymbol{X}_{t}$ can lead to misleading estimation, resulting in a wrong model fit.
@@
In particular, we consider $\boldsymbol{X}_{t}\sim\operatorname{NPN}(0,H_{t},f)$ where $\operatorname{NPN}$ is the nonparanormal distribution (liu2009nonparanormal) and the set of functions $\{f_{j}\}_{j=1}^{p}$ transform the the random vector $\boldsymbol{X}_{t}=\left(X_{1,t},\ldots,X_{p,t}\right)^{\prime}$, into the random vector $f(\boldsymbol{X}_{t})=\left(f_{1}\left(X_{t,1}\right),\ldots,f_{p}\left(X_{p,t}\right)\right)^{\prime}$, which follows a multivariate Gaussian distribution. The nonparanormal distribution depends on the transformation functions $\{f_{j}\}$ and the covariance matrix $H_{t}$, both estimated from the data.
@@
When data are truly normally distributed, Spearman’s rho and Kendall’s tau statistics may be used to determine the unknown correlation matrix exploiting their relation with the Pearson correlation coefficient (kruskal1958ordinal). This approach is suggested in the SKEPTIC [^1] Spearman’s rho and Kendall’s tau estimates preempt to infer correlation. estimator (liu2012high), a nonparametric and robust estimator of the correlation matrix that is less sensitive to outliers and does not strictly rely on the normality assumption as the Pearson correlation coefficient does.
@@
This paper adds to two main research fields: (i) the modeling of temporal dependence using a copula function (beare2010copulas, patton2012review, patton2013copula and fan2014copulas), and (ii) the nonlinear modeling of multivariate time series with a copula-based function, (chen2006estimation, lee2009copula, remillard2012copula, fan2023estimation and cordoni2024consistent). In particular, fan2023estimation consider a semiparametric Gaussian copula process for estimating the transition matrix within a VAR model, whereas cordoni2024consistent use it for causal inference in a high-dimensional time series context.
@@ 2.2 Dynamic Conditional Corelation
To model $R_{t}$, engle2002dynamic assumes that $\boldsymbol{X}_{t}\sim N(0,H_{t})$, where the residuals $\varepsilon_{t}=D_{t}^{-1}\boldsymbol{X}_{t}$ are such that $\mathbb{E}\left(\varepsilon_{t}\varepsilon_{t}^{\prime}\mid\mathcal{F}_{t-1}\right)=R_{t}$ is the unconditional correlation matrix of the returns at time $t$, modeled according to the following linear transformation:
@@
In large dimensions, the estimation of the unconditional correlation matrix $\bar{Q}$ suffers from overfitting due to the curse of the dimensionality problem. engle2019large solve this issue with the DCC-NL model, introducing a non-linear shrinkage of the eigenvalues, $\bar{Q}$, using the QuEST function in ledoit2015spectrum that produces a more stable estimate of the matrix.
@@ 3.1 Nonparanormal
where $f(x_{t})=D_{t}f(\varepsilon_{t})$ and $\varepsilon_{t}=D_{t}^{-1}x_{t}$, so that the same conditional scale matrix $D_{t}$ governs both the standardization and the subsequent transformation. so that $f(x_{t})$ applies the monotone transformations to the standardized components of $\boldsymbol{X}_{t}$. We therefore write $\boldsymbol{X}_{t}\sim\operatorname{NPN}(0,H_{t},f).$
@@
where $\Phi_{R_{t}}$ denotes the multivariate Gaussian cumulative distribution function (CDF) with correlation matrix $R_{t}$.
@@
fan2023estimation and cordoni2024consistent extend the nonparanormal framework to dependent data, proposing vector autoregressive structures for latent Gaussian scores. In our setting, we focus instead on modeling the second-order dynamics through $R_{t}$, leading to the Dynamic Conditional SKEPTIC specification.
@@ 3.2 Dynamic Conditional SKEPTIC
where $\zeta_{is}$ and $\zeta_{js}$ are the ranks of $X_{is}$ and $X_{js}$ among the observations $(X_{i1},\dots,X_{iT})$ and $(X_{j1},\dots,X_{jT})$, respectively, and $\bar{\zeta}_{i}=\bar{\zeta}_{j}=(T+1)/2$ denotes the average rank. For a generic time $t$, these statistics provide nonparametric estimates of the dependence between the empirical realizations of $X_{it}$ and $X_{jt}$. Let $\widetilde{X}_{it}$ and $\widetilde{X}_{jt}$ be independent copies of $X_{it}$ and $X_{jt}$, and let $\Phi_{i}$ and $\Phi_{j}$ denote their marginal distribution functions. Then, the population versions of Spearman’s rho and Kendall’s tau are defined as: $\rho^{s}_{ij}:=\operatorname{Corr}\left(\Phi_{i}(X_{it}),\Phi_{j}(X_{jt})\right)$, and $\tau_{ij}:=\operatorname{Corr}\left(\operatorname{sign}(X_{it}-\widetilde{X}_{it}),\operatorname{sign}(X_{jt}-\widetilde{X}_{jt})\right).$ Substituting $\widehat{R}^{\rho}_{t}$ or $\widehat{R}^{\tau}_{t}$ to the canonical correlation matrix used in engle2002dynamic, we introduce the Dynamic Conditional SKEPTIC.
@@ 4 Theoretical Properties
We establish the theoretical foundation of the SKEPTIC estimator for an autoregressive process, demonstrating that this novel approach is a reliable extension to the DCC model. We start recalling the stationarity conditions outlined in the work of fermanian2017stationarity, where the DCC is reformulated in terms of a Markov Chain. In our case, under the assumption of nonparanormal distribution, we define the following model:
@@
for any $t$, where $T:=[T_{ij,t}]_{1\leq i,j\leq 4}$ is a block matrix. [^2] We refer to fermanian2017stationarity, where $\mu=\nu=r=s=1$. To determine the stationarity of the DCS process, we define the data-generating process of $W_{t}$:
@@
(Stationarity of DCC (fermanian2017stationarity)) Given the process $\delta_{t}$ in (11), we state that:
@@ 4.1 Mixing Conditions of Semiparametric Gaussian Process
(beare2010copulas) The $\beta$-mixing coefficients $\left\{\beta_{k}:k\in\mathbb{N}\right\}$ that correspond to the sequence of random variables $\left\{X_{t}\right\}$ are given by
@@
The proof is in the Appendix A. Moreover, by Proposition 2, and 4 in remillard2012copula, the Gaussian copula process is also a $\rho$-mixing process, where the $\rho$-mixing coefficients are defined as follows.
@@
(beare2010copulas) The $\rho$-mixing coefficients $\left\{\rho_{k}:k\in\mathbb{N}\right\}$ that correspond to the sequence of random variables $\left\{X_{t}\right\}$ are given by
@@
The semiparametric Gaussian copula exhibits strong mixing properties, characterizing a process where serial dependence among random variables diminishes over time. We exploit these mixing conditions to derive concentration inequalities for the Dynamic Conditional SKEPTIC. We recall that Spearman’s rho and Kendall’s tau can be expressed in terms of U-statistics. The U-statistic of Kendall’s tau is:
@@ 5 Estimation
$L_{C}(\theta,\phi)=\frac{1}{2}\sum_{t=1}^{T}\left(\log\left|\mathcal{R}_{t}\right|+\nu_{t}^{\prime}R_{t}^{-1}\nu_{t}+\nu_{t}^{\prime}\nu_{t}\right).$ (15)
@@
The estimation of $\phi=(\alpha,\beta)$ proposed in engle2002dynamic leads to major issues as dimensionality grows. This difficulty aligns with the curse of dimensionality described in engle2019large. Specifically, as the concentration ratio $p/T$ increases, the estimation of $\mathcal{R}_{t}$ suffers from overfitting, and when $p/T$ is equal to 1, the matrix $\mathcal{R}_{t}$ becomes singular and the procedure inapplicable.
We follow pakel2021fitting and use the composite likelihood function that estimates the parameters by summing the quasi-likelihoods of asset subsets, thereby addressing lower-dimensional marginal densities. This method involves determining the conditional variances with a GARCH-type model and then fitting a likelihood based on correlation $2\times 2$. More precisely, once the values of $\hat{\theta}$ that maximize (14) are obtained (fermanian2024model), we estimate the parameter set $\phi$ maximizing:
@@ 5.1 Asymptotic Properties of Composite Likelihood Estimator
pakel2021fitting demonstrate that the composite likelihood estimator yields consistent and asymptotically normal estimates.
@@ 6 Simulation Studies
We explore the properties of the Dynamic Conditional SKEPTIC by conducting Monte Carlo simulations. To make the simulations realistic, we first estimate the unconditional covariance matrix of the S&P 500 constituents, denoted by $H_{t}$. We then generate a sample of 1000 simulated returns according to the following dynamics. Let $\boldsymbol{X}_{t}$ denote the vector of returns; we assume
$x_{t}=H^{1/2}_{t}u_{t},$ (20)
where $u_{t}\sim t_{\nu}\left(0,\sigma_{t}\right)$, and the number of degrees of freedom $\nu$ is randomly drawn between 4 and 8. As is well known, empirical returns exhibit non-negligible skewness. To reproduce this feature, we introduce contamination into the simulated data. Let $\delta$ denote the contamination level, we consider three cases, $\delta\in$ {0.01,0.05,0.100.01,0.05,0.10}, corresponding to $1\%,5\%$, and $10\%$ of contaminated observations, respectively. The contamination is introduced by replacing the specified proportion of the data with draws from a $t_{\nu}\left(0,\sigma_{t}\right)$ distribution with $\nu=3$, thereby increasing the number of outliers and inducing asymmetry in the simulated return distribution. We then simulate a DCC model with parameters $\alpha=0.02$ and $\beta=0.97$ as in pakel2021fitting.
For each level of $\delta$, we estimate the parameters of the DCC-based processes and the DCS models $\{\alpha,\beta\}$. We perform the simulations 100 times. We consider four different dimensions: $p\in\{50,97,250,429\}$, where $p=429$, and $p=97$ correspond respectively to the dimension of the S&P500 and S&P100 constituents. In the cases of $p=50$, and $p=250$ we randomly select stocks from the constituents of S&P500 to estimate $H_{t}$. The number of observations we simulate are: $T\in\{500,1000,2000\}$. The results of the simulation are shown in Appendix C. The simulation results highlight that the persistence parameter $\beta$ in the DCS models is lower than the DCC-based specification. In contrast, the $\alpha$ parameters are average higher in the DCS models and closer to the true values than those obtained from the DCC model. The larger $\alpha$ coefficients suggest that the DCS model responds more promptly to short-term fluctuations in dependence. Conversely, the DCC and DCC-NL models, which rely on Gaussian assumptions, tend to smooth out such variations and exhibit a slower adjustment to shocks.
It is worth noting to highlight that as the level of contamination, explained by $\delta$ increases, the standard deviation of the estimates of the DCS model is lower than the DCC-based models, highlighting the ability of the Dynamic Conditional SKEPTIC to mitigate the impact of outlier values.
@@ 7 Empirical Analysis
We conduct an empirical analysis comparing the Dynamic Conditional SKEPTIC (DCS) with the DCC model of engle2002dynamic and the DCC-NL of engle2019large [^3] We thank Michael Wolf for kindly providing the codes for the DCC and DCC-NL on the website https://www.econ.uzh.ch/en/people/faculty/wolf/publications.html under the link “Programming Code.” We revisit these codes, aligning them to our purposes.. We download stock price data from Yahoo! Finance of the stocks that belong to the S&P100 and S&P500 indices, covering the period from 02/01/2013 to 23/01/2025. After cleaning the data, the final dataset includes 429 stocks for the S&P500 and 97 stocks for the S&P100, each with 3034 price observations.

### v2

@@ 1 Introduction
Multivariate GARCH models analyze the dependence structure of a vector of asset returns through the dynamic estimation of its covariance matrix. When the number of assets is large, these models face computational issues in the estimation procedure that returns poor performance due to the curse of dimensionality. To this end, the recent works of Engle et al. [10] and De Nard et al. [8] robustify the Dynamic Conditional Correlation (DCC) (12), one of the most commonly used in the class of multivariate GARCH models, for estimating time-varying correlations in financial assets, including a nonlinear shrinkage (22) that improves the estimation of the unconditional correlation matrix and relying on the composite likelihood method (28) that provides computationally feasible estimation of the DCC model in large-dimensional settings.
However, a limitation of the DCC model remains the assumption that asset returns follow a normal distribution. This assumption does not adequately capture the empirical stylized facts [6] which are commonly observed in asset return distributions [2], particularly the skewness and leptokurtosis. To formalize the concept, let $\boldsymbol{X}_{t}=(X_{1,t},\ldots,X_{p,t})^{\prime}$ represent the $p$-dimensional random vector of asset returns, the main assumption is that $\boldsymbol{X}_{t}\sim\operatorname{N}(0,H_{t})$, where $H_{t}$ is the covariance matrix. The difference between the normal distribution and the empirical distribution of $\boldsymbol{X}_{t}$ can lead to misleading estimation, resulting in a wrong model fit.
@@
In particular, we consider $\boldsymbol{X}_{t}\sim\operatorname{NPN}(0,H_{t},f)$ where $\operatorname{NPN}$ is the nonparanormal distribution (25) and the set of functions $\{f_{j}\}_{j=1}^{p}$ transform the the random vector $\boldsymbol{X}_{t}=\left(X_{1,t},\ldots,X_{p,t}\right)^{\prime}$, into the random vector $f(\boldsymbol{X}_{t})=\left(f_{1}\left(X_{t,1}\right),\ldots,f_{p}\left(X_{p,t}\right)\right)^{\prime}$, which follows a multivariate Gaussian distribution. The nonparanormal distribution depends on the transformation functions $\{f_{j}\}$ and the covariance matrix $H_{t}$, both estimated from the data.
@@
When data are truly normally distributed, Spearman’s rho and Kendall’s tau statistics may be used to determine the unknown correlation matrix exploiting their relation with the Pearson correlation coefficient (20). This approach is suggested in the SKEPTIC [^1] Spearman’s rho and Kendall’s tau estimates preempt to infer correlation. estimator (24), a nonparametric and robust estimator of the correlation matrix that is less sensitive to outliers and does not strictly rely on the normality assumption as the Pearson correlation coefficient does.
@@
This paper adds to two main research fields: (i) the modeling of temporal dependence using a copula function (3, 29, 30 and 14), and (ii) the nonlinear modeling of multivariate time series with a copula-based function, (4, 23, 31, 13 and 7). In particular, Fan et al. [13] consider a semiparametric Gaussian copula process for estimating the transition matrix within a VAR model, whereas Cordoni and Sancetta [7] use it for causal inference in a high-dimensional time series context.
@@ 2.2 Dynamic Conditional Corelation
To model $R_{t}$, Engle [12] assumes that $\boldsymbol{X}_{t}\sim N(0,H_{t})$, where the residuals $\varepsilon_{t}=D_{t}^{-1}\boldsymbol{X}_{t}$ are such that $\mathbb{E}\left(\varepsilon_{t}\varepsilon_{t}^{\prime}\mid\mathcal{F}_{t-1}\right)=R_{t}$ is the unconditional correlation matrix of the returns at time $t$, modeled according to the following linear transformation:
@@
In large dimensions, the estimation of the unconditional correlation matrix $\bar{Q}$ suffers from overfitting due to the curse of the dimensionality problem. Engle et al. [10] solve this issue with the DCC-NL model, introducing a non-linear shrinkage of the eigenvalues, $\bar{Q}$, using the QuEST function in Ledoit and Wolf [22] that produces a more stable estimate of the matrix.
@@ 3.1 Nonparanormal
where $f(x_{t})=D_{t}f(\varepsilon_{t})$, so that the same conditional scale matrix $D_{t}$ governs both the standardization and the subsequent transformation, so that $f(x_{t})$ applies the monotone transformations to the standardized components of $\boldsymbol{X}_{t}$. We therefore write $\boldsymbol{X}_{t}\sim\operatorname{NPN}(0,H_{t},f).$
@@
where $\Phi_{R_{t}}$ denotes the multivariate Gaussian cumulative distribution function (CDF) with correlation matrix $R_{t}$, and $\Phi^{-1}(\cdot)$ denotes the univariate quantile function of a standard normal distribution.
@@
[13] and [7] extend the nonparanormal framework to dependent data, proposing vector autoregressive structures for latent Gaussian scores. In our setting, we focus instead on modeling the second-order dynamics through $R_{t}$, leading to the Dynamic Conditional SKEPTIC specification.
@@ 3.2 Dynamic Conditional SKEPTIC
where $\zeta_{is}$ and $\zeta_{js}$ are the ranks of $X_{is}$ and $X_{js}$ among the observations $(X_{i1},\dots,X_{iT})$ and $(X_{j1},\dots,X_{jT})$, respectively, and $\bar{\zeta}_{i}=\bar{\zeta}_{j}=(T+1)/2$ denotes the average rank. For a generic time $t$, these statistics provide nonparametric estimates of the dependence between the empirical realizations of $X_{it}$ and $X_{jt}$. Let $\widetilde{X}_{it}$ and $\widetilde{X}_{jt}$ be independent copies of $X_{it}$ and $X_{jt}$, and let $\Phi_{i}$ and $\Phi_{j}$ denote their marginal distribution functions. Then, the population versions of Spearman’s rho and Kendall’s tau are defined as: $\rho^{s}_{ij}:=\operatorname{Corr}\left(\Phi_{i}(X_{it}),\Phi_{j}(X_{jt})\right)$, and $\tau_{ij}:=\operatorname{Corr}\left(\operatorname{sign}(X_{it}-\widetilde{X}_{it}),\operatorname{sign}(X_{jt}-\widetilde{X}_{jt})\right).$ Substituting $\widehat{R}^{\rho}_{t}$ or $\widehat{R}^{\tau}_{t}$ to the canonical correlation matrix used in Engle [12], we introduce the Dynamic Conditional SKEPTIC.
@@ 4 Theoretical Properties
We establish the theoretical foundation of the SKEPTIC estimator for an autoregressive process, demonstrating that this novel approach is a reliable extension to the DCC model. We start recalling the stationarity conditions outlined in the work of Fermanian and Malongo [15], where the DCC is reformulated in terms of a Markov Chain. In our case, under the assumption of nonparanormal distribution, we define the following model:
@@
for any $t$, where $T:=[T_{ij,t}]_{1\leq i,j\leq 4}$ is a block matrix. [^2] We refer to Fermanian and Malongo [15], where $\mu=\nu=r=s=1$. To determine the stationarity of the DCS process, we define the data-generating process of $W_{t}$:
@@
(Stationarity of DCC (15)) Given the process $\delta_{t}$ in (11), we state that:
@@ 4.1 Mixing Conditions of Semiparametric Gaussian Process
(3) The $\beta$-mixing coefficients $\left\{\beta_{k}:k\in\mathbb{N}\right\}$ that correspond to the sequence of random variables $\left\{X_{t}\right\}$ are given by
@@
The proof is in the Appendix A. Moreover, by Proposition 2, and 4 in Rémillard et al. [31], the Gaussian copula process is also a $\rho$-mixing process, where the $\rho$-mixing coefficients are defined as follows.
@@
(3) The $\rho$-mixing coefficients $\left\{\rho_{k}:k\in\mathbb{N}\right\}$ that correspond to the sequence of random variables $\left\{X_{t}\right\}$ are given by
@@
The semiparametric Gaussian copula process exhibits strong mixing properties, characterizing a process where serial dependence among random variables diminishes over time. We exploit these mixing conditions to derive concentration inequalities for the Dynamic Conditional SKEPTIC. We recall that Spearman’s rho and Kendall’s tau can be expressed in terms of U-statistics. The U-statistic of Kendall’s tau is:
@@ 5 Estimation
$L_{C}(\theta,\phi)=\frac{1}{2}\sum_{t=1}^{T}\left(\log\left|\mathcal{R}_{t}\right|+\nu_{t}^{\prime}\mathcal{R}_{t}^{-1}\nu_{t}+\nu_{t}^{\prime}\nu_{t}\right).$ (15)
@@
The estimation of $\phi=(\alpha,\beta)$ proposed in Engle [12] leads to major issues as dimensionality grows. This difficulty aligns with the curse of dimensionality described in Engle et al. [10]. Specifically, as the concentration ratio $p/T$ increases, the estimation of $\mathcal{R}_{t}$ suffers from overfitting, and when $p/T$ is equal to 1, the matrix $\mathcal{R}_{t}$ becomes singular and the procedure inapplicable.
We follow Pakel et al. [28] and use the composite likelihood function that estimates the parameters by summing the quasi-likelihoods of asset subsets, thereby addressing lower-dimensional marginal densities. This method involves determining the conditional variances with a GARCH-type model and then fitting a likelihood based on correlation $2\times 2$. More precisely, once the values of $\hat{\theta}$ that maximize (14) are obtained (16), we estimate the parameter set $\phi$ maximizing:
@@ 5.1 Asymptotic Properties of Composite Likelihood Estimator
Pakel et al. [28] demonstrate that the composite likelihood estimator yields consistent and asymptotically normal estimates.
@@ 6 Simulation Studies
We assess the ability of the Dynamic Conditional SKEPTIC to recover the true dependence dynamics in finite samples by comparing its parameter estimates with those obtained from DCC and DCC-NL models in a Monte Carlo setting.
We first compute an empirical target matrix $\bar{Q}$ from log-returns of the selected stocks, which is used as the intercept in the correlation recursion. For each Monte Carlo replication, we simulate a $p$-dimensional return series $\{x_{t}\}_{t=1}^{T}$ according to the DCC model in (6), with fixed parameters $\alpha=0.02,\beta=0.97$. Recalling that $H_{t}=D_{t}R_{t}D_{t}$, we generate returns as $x_{t}=H_{t}^{1/2}u_{t}$, $\forall t\in\{1,\dots,T\}$. For each asset $i=1,\ldots,p$, the innovation $u_{i,t}$ is drawn from a Student-$t$ distribution $t_{\nu}\!\left(0,1\right)$, and the number of degrees of freedom $\nu$ is randomly drawn between 4 and 8.
The empirical returns exhibit non-negligible skewness. To reproduce this feature, we introduce contamination into the simulated data. Let $\delta$ denote the contamination level; we consider three cases, $\delta\in\{0.01,0.05,0.10\}$, corresponding to $1\%$, $5\%$, and $10\%$ of contaminated observations, respectively. The contamination is introduced by replacing, independently across assets and time, the specified proportion of the innovations $u_{i,t}$ with draws from a $t_{\nu}\!\left(0,1\right)$ distribution with $\nu=3$, thereby increasing the number of outliers and inducing asymmetry in the simulated return distribution.
[…]
@@ 7 Empirical Analysis
We conduct an empirical analysis comparing the Dynamic Conditional SKEPTIC (DCS) with the DCC model of Engle [12] and the DCC-NL of Engle et al. [10] [^3] We thank Michael Wolf for kindly providing the codes for the DCC and DCC-NL on the website https://www.econ.uzh.ch/en/people/faculty/wolf/publications.html under the link “Programming Code.” We revisit these codes, aligning them to our purposes.. We download stock price data from Yahoo! Finance of the stocks that belong to the S&P100 and S&P500 indices, covering the period from 02/01/2013 to 23/01/2025. After cleaning the data, the final dataset includes 429 stocks for the S&P500 and 97 stocks for the S&P100, each with 3034 price observations.